OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2 -5 5 9 -9 when I tried I got 8

OpenStudy (anonymous):

@ivettef365

OpenStudy (compassionate):

-4(2x - 1) = -6(x + 2) - 2 -8x + 4 = -6x - 12 - 2 -8x + 4 = -6x - 14 -8x = -6x - 16 -2x = -16 x = 8. You are correct.

OpenStudy (anonymous):

yeah but the answer choices are -5, 5, 9 or -9

OpenStudy (compassionate):

Read it again.

OpenStudy (anonymous):

Yeah but 8 isn't an answer choice.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (compassionate):

x = 9.

OpenStudy (anonymous):

thanks